If a sample of size 38 is collected from finite population of 600, then the finite population correction factor is equal to


1
Expert's answer
2021-03-03T04:18:12-0500

Let finite population correction be FPC


"pfc=\sqrt{\frac{N-n}{N-1}}"


N= 600 and n = 38


"pfc=\sqrt{\frac{600-38}{600-1}}"


"pfc=\sqrt{\frac{562}{599}}"


"fpc = 0.9686"